6c+5d-4c-3d+3c-6d when c=4 and d=(-2)
10a+3b-5a+4b+1a+5b when a=(-3) and b=5
3m+9n+6m-7n-4m+2n when m=6 and n=(-4)

Answer by checkley77(12844)

6c+5d-4c-3d+3c-6d when c=4 and d=(-2)
6*4+5*-2-4*4-3*-2+3*4-6*-2
24-10-16+6+12+12
54-26=28 ans.
-----------------------------
10a+3b-5a+4b+1a+5b when a=(-3) and b=5
10*-3+3*5-5*-3+4*5+1*-3+5*5
-30+15+15+20-3+25
75-33=42 ans.
------------------------------
I'll leave this last one for you to practice with.
Just substitute 6 for m & -4 for n.
Multiply then add & subtract.
------------------------------
3m+9n+6m-7n-4m+2n when m=6 and n=(-4)